what is f(t)=1/2t+2 when t=2,4,8
Answer by Alan3354(69443)

what is f(t)=1/2t+2 when t=2,4,8
If you mean (1/2)t + 2
------------------------
f(2) = (1/2)*2 + 2 = 1 + 2 = 3
-----------
Do 4 and 8 the same way.